# Hierarchy-Boosted Funnel Learning for Identifying Semiconductors with Ultralow Lattice Thermal Conductivity

### v1

@@ II.5 Mechanisms of Phonon Thermal Transport Properties
Generally speaking, compounds with heavier atoms tend to exhibit lower group velocities due to the reduced phonon frequency. As a result, a lower $\kappa_{\mathrm{L}}$ of Cs2SnSe3 was expected given the relatively heavy nature of Sn. However, it is noteworthy that the $\kappa_{\mathrm{L}}$ of Cs2SnSe3 is obviously higher that that of Cs2GeSe3 (i.e., $\sim$3 times in the a-axis, $\sim$1.7 times in the b-axis, $\sim$2 times in the c-axis at 300 K), which presents an interesting unusual phenomenon. Next, we would like to discuss the microscopic mechanisms responsible for the ultralow $\kappa_{\mathrm{L}}$ and unusual difference observed among the two materials.
@@
The phonon dispersion reveals that there are 3 acoustic branches and 33 optical branches at each phonon wave vector q due to the 12 atoms in the primitive cell of both Cs2SnSe3 and Cs2GeSe3, in which three acoustic branches are composed of one longitudinal acoustic mode (LA) and two transverse acoustic modes (TA and TA’). A striking common feature of their phonon dispersion is that a waterfall-like low-lying optical branch (LLO) exhibits the avoided crossing behavior with acoustic modes around the Brillouin zone center, resulting in strong acoustic-optical coupling as one of the potential signals of the rattling model.[69] These characteristics can not only lead to a softening of the acoustic modes thereby yielding low phonon group velocities, but also greatly enhance the scattering rates of heat-carrying acoustic phonons, all of which contribute to the suppression of $\kappa_{\mathrm{L}}$ for Cs2SnSe3 and Cs2GeSe3. Analysis of the atom-projected PDOS and spectral $\kappa_{\mathrm{L}}(\omega)$ (middle and right panels in Figures 5c and 5d) indicate that acoustic and LLO phonon modes in the low-frequency range (0–2 THz) are primarily dictated by Cs and Se vibrations followed by Sn/Ge vibrations, which also make predominant contributions to the ultralow $\kappa_{\mathrm{L}}$. A localized region mostly contributed by Cs atoms within a narrow energy window centered around 1 THz is observed in both compounds, indicating the anharmonic rattling-like motion of the weakly bonded Cs atoms, which is responsible for the presence of the soft LLO modes. This can also be confirmed in the potential energy curves obtained by shifting the atoms with respect to their static equilibrium positions along different axes as depicted in Figure S3. Sn/Ge and Se atoms are both confined within the comparatively steep potential wells, whereas Cs atoms with heavier mass can vibrate easily with larger amplitude due to the shallowest potential energy surface in all the direction. These atoms exhibit nearly identical displacement magnitudes in their respective compounds. Therefore, the loosely bound Cs atoms are surrounded by [Sn2Se6]/[Ge2Se6] units thereby exhibiting the same rattling-like effect, which can result in a common strong anharmonicity in both compounds.[70, 71]
@@
Figures 5e and S4 show the frequency dependence of the group velocity $\upsilon$ from all the q points in different axes at 300 K for Cs2SnSe3 and Cs2GeSe3, as given by $\upsilon_{\lambda}={\partial\omega_{\lambda}}/{\partial q}$. Most of the phonon modes possess ultralow $\upsilon$ values less than 2 km/s, confirming the lattice softening induced by the weak interatomic bonding, which is also consistent with the flat phonon bands observed in the phonon dispersion. Figure 5f depicts the overall short $\tau$ values mainly ranging from 0.2 ps to 20 ps, with low-frequency phonons exhibiting a gradual shortening trend in both compounds. This can be attributed to the presence of LLO modes, which facilitates more scattering paths and impedes the heat flow, significantly enhancing the scattering rates thereby reducing $\tau$ for both acoustic and LLO modes, ultimately reducing $\kappa_{\mathrm{L}}$. It is noteworthy that in the low-frequency phonon region, particularly within the acoustic phonon range, the phonon group velocities of these two compounds show little difference, yet the phonon lifetime of Cs2SnSe3 is significantly longer than that of Cs2GeSe3. Therefore, we conclude that the relatively lower anomalous $\kappa_{\mathrm{L}}$ of Cs2GeSe3 is attributed to its shorter phonon lifetime.
@@
We speculate that the difference in the phonon lifetime between Cs2SnSe3 and Cs2GeSe3 might be related to the strength of Sn–Se and Ge–Se covalent bonding. This covalent bonding may induce anisotropic motion, involving the collective movement of Sn/Ge and Se atoms, thereby leading to a strong lattice anharmonicity within the system.[72] To support this conclusion, we perform crystal orbital Hamilton population (COHP) calculations[73] to identify the energy-resolved local bonding information. Figure 5g shows that antibonding states persist down to –3 eV below the Fermi level ($E$F) in Sn–Se bonding, which can weaken the corresponding bonding strength thereby forming a polar covalent bond as observed in ELF. This can be demonstrated by the integral of COHP ($\mathrm{ICOHP}=\int_{-\infty}^{E}\mathrm{COHP}(E)\mathrm{d}E$) at the $E$F, which represents all occupied orbitals, serving as an indicator of bond strength. Since the average ICOHP values for Sn–Se and Ge–Se bonds are –0.72 eV and –4.49 eV, respectively, indicating the obvious weaker Sn–Se bonding in Cs2SnSe3. The stronger bonding strength of Ge–Se is also evident from its larger force constant $\left|\Phi_{ij}\right|$ in comparison to that of Sn–Se (Figure S5). Hence, the shorter phonon lifetime of Cs2GeSe3 relative to Cs2SnSe3 may be attributed to their significant difference in Ge–Se and Sn–Se covalent bonding, which can be further quantified at the structural descriptor level through subsequent interpretable ML approaches.
@@ III Conclusion
In summary, we propose a novel HiBoFL framework via integrating unsupervised learning and supervised learning to efficiently predict complex properties and uncover structure-property relationships. As a compelling demonstration, this framework has been applied to efficiently identify semiconductors with ultralow $\kappa_{\mathrm{L}}$, which circumvents large-scale brute-force DFT calculations without clear objectives. By employing unsupervised learning for the materials from the MP database, we successfully group them into seven clusters. A few hundred materials in clusters C1 and C2 with a high likelihood of possessing low $\kappa_{\mathrm{L}}$ is selected from a pool of hundreds of thousands based on similarity design rules. We further conduct low-cost HTC on materials belonging to the two clusters, establishing a local database for researchers to retrieve and providing a list of candidate materials with potential applications in the TE field. Additionally, statistical analysis of these candidates offers valuable domain knowledge to guide the design of materials with ultralow $\kappa_{\mathrm{L}}$. Cs2GeSe3 and Cs2SnSe3 with ultralow $\kappa_{\mathrm{L}}$ ($\sim$0.25 W/mK) are screened out, in which the anomalous ultralow $\kappa_{\mathrm{L}}$ of Cs2GeSe3 is attributed to the lower $\tau$ caused by the difference in covalent bonding. Based on the established local database, we train a robust ML classification model to refine unsupervised learning, achieving a ROC AUC of 0.94 on the test set, which can enable the efficient predictions of ultralow $\kappa_{\mathrm{L}}$ materials. The interpretable analysis of the classification model reveals the mechanisms of key descriptors on $\kappa_{\mathrm{L}}$ modulation, such as $W_{\mathrm{A}}^{\mathrm{min}}$, $\alpha$ and $L^{\mathrm{min}}$, etc. The special factor $L^{\mathrm{min}}$ is discovered to show a unique influence on structural anharmonicity, leading to the difference of phonon lifetime in Cs2GeSe3 and Cs2SnSe3, in agreement with the first-principles analysis. We believe that this work provides a novel feasible way for efficiently seeking promising TE materials, in which the proposed HiBoFL framework is also expected to be applied in other material fields.
@@ IV.2 Theoretical Framework
Where $\overline{\upsilon}_{\mathrm{l}}$, $\overline{\upsilon}_{\mathrm{t}}$, $\overline{\gamma}_{\mathrm{l}}$ and $\overline{\gamma}_{\mathrm{t}}$ are the longitudinal sound velocity, transverse sound velocity, longitudinal Grüneisen parameter and transverse Grüneisen parameter, respectively. B and G are bulk moduli and shear moduli, respectively. In this manner, we can calculate the $\kappa_{\mathrm{PET}}$ for different materials derived from their elastic properties within the HTC framework. To further validate the $\kappa_{\mathrm{PET}}$ from the PET empirical equation, we calculated more accurate $\kappa_{\mathrm{L}}$ values for Cs2SnSe3 and Cs2SnSe3 by solving the phonon Boltzmann transport equation.
IV.3 Machine Learning Toolkit
The chemical compositions and crystal structures were featurized into different descriptors using the Matminer package.[86] All parts related to ML were carried out using the Scikit-learn package.[87] To achieve the automation and acceleration of hyperparameter optimization in supervised learning, we employed the powerful Optuna package[88] for efficiently finding the best hyperparameters. And to provide the interpretable analysis of the black-box ML model, a game theoretic approach as implemented in the SHAP package[45, 46] was employed.

### v2

@@ II.5 Mechanisms of Phonon Thermal Transport Properties
Generally speaking, compounds with heavier atoms tend to exhibit lower group velocities due to the reduced phonon frequency. As a result, a lower $\kappa_{\mathrm{L}}$ of Cs2SnSe3 was expected given the relatively heavy nature of Sn. However, it is noteworthy that the $\kappa_{\mathrm{L}}$ of Cs2SnSe3 is obviously higher than that of Cs2GeSe3 (i.e., $\sim$3 times in the a-axis, $\sim$1.7 times in the b-axis, $\sim$2 times in the c-axis at 300 K), which presents an interesting unusual phenomenon. Next, we would like to discuss the microscopic mechanisms responsible for the ultralow $\kappa_{\mathrm{L}}$ and unusual difference observed among the two materials.
@@
The phonon dispersion reveals that there are 3 acoustic branches and 33 optical branches at each phonon wave vector q due to the 12 atoms in the primitive cell of both Cs2SnSe3 and Cs2GeSe3, in which three acoustic branches are composed of one longitudinal acoustic mode (LA) and two transverse acoustic modes (TA and TA’). A striking common feature of their phonon dispersion is that a waterfall-like low-lying optical branch (LLO) exhibits the avoided crossing behavior with acoustic modes around the Brillouin zone center, resulting in strong acoustic-optical coupling as one of the potential signals of the rattling model.[74] These characteristics can not only lead to a softening of the acoustic modes thereby yielding low phonon group velocities, but also greatly enhance the scattering rates of heat-carrying acoustic phonons, all of which contribute to the suppression of $\kappa_{\mathrm{L}}$ for Cs2SnSe3 and Cs2GeSe3. Analysis of the atom-projected PDOS and spectral $\kappa_{\mathrm{L}}(\omega)$ (middle and right panels in Figures 5c and 5d) indicate that acoustic and LLO phonon modes in the low-frequency range (0–2 THz) are primarily dictated by Cs and Se vibrations followed by Sn/Ge vibrations, which also make predominant contributions to the ultralow $\kappa_{\mathrm{L}}$. A localized region mostly contributed by Cs atoms within a narrow energy window centered around 1 THz is observed in both compounds, indicating the anharmonic rattling-like motion of the weakly bonded Cs atoms, which is responsible for the presence of the soft LLO modes. This can also be confirmed in the potential energy curves obtained by shifting the atoms with respect to their static equilibrium positions along different axes as depicted in Figure S3. Sn/Ge and Se atoms are both confined within the comparatively steep potential wells, whereas Cs atoms with heavier mass can vibrate easily with larger amplitude due to the shallowest potential energy surface in all the directions. These atoms exhibit nearly identical displacement magnitudes in their respective compounds. Therefore, the loosely bound Cs atoms are surrounded by [Sn2Se6]/[Ge2Se6] units thereby exhibiting the same rattling-like effect, which can result in a common strong anharmonicity in both compounds.[75, 76]
@@
Figures 5e and S4 show the frequency dependence of the group velocity $\upsilon$ from all the q points in different axes at 300 K for Cs2SnSe3 and Cs2GeSe3, as given by $\upsilon_{\lambda}={\partial\omega_{\lambda}}/{\partial q}$. Most of the phonon modes possess ultralow $\upsilon$ values less than 2 km/s, confirming the lattice softening induced by the weak interatomic bonding, which is also consistent with the flat phonon bands observed in the phonon dispersion. Figure 5f depicts the overall short $\tau$ values mainly ranging from 0.2 ps to 20 ps, with low-frequency phonons exhibiting a gradual shortening trend in both compounds. This can be attributed to the presence of LLO modes, which facilitate more scattering paths and impede the heat flow, significantly enhancing the scattering rates thereby reducing $\tau$ for both acoustic and LLO modes, ultimately reducing $\kappa_{\mathrm{L}}$. It is noteworthy that in the low-frequency phonon region, particularly within the acoustic phonon range, the phonon group velocities of these two compounds show little difference, yet the phonon lifetime of Cs2SnSe3 is significantly longer than that of Cs2GeSe3. Therefore, we conclude that the relatively lower anomalous $\kappa_{\mathrm{L}}$ of Cs2GeSe3 is attributed to its shorter phonon lifetime.
@@
We speculate that the difference in the phonon lifetime between Cs2SnSe3 and Cs2GeSe3 might be related to the strength of Sn–Se and Ge–Se covalent bonding. This covalent bonding may induce anisotropic motion, involving the collective movement of Sn/Ge and Se atoms, thereby leading to a strong lattice anharmonicity within the system.[77] To support this conclusion, we perform crystal orbital Hamilton population (COHP) calculations[78] to identify the energy-resolved local bonding information. Figure 5g shows that anti-bonding states persist down to –3 eV below the Fermi level ($E$F) in Sn–Se bonding, which can weaken the corresponding bonding strength thereby forming a polar covalent bond as observed in ELF. This can be demonstrated by the integral of COHP ($\mathrm{ICOHP}=\int_{-\infty}^{E}\mathrm{COHP}(E)\mathrm{d}E$) at the $E$F, which represents all occupied orbitals, serving as an indicator of bond strength. Since the average ICOHP values for Sn–Se and Ge–Se bonds are –0.72 eV and –4.49 eV, respectively, indicating the obviously weaker Sn–Se bonding in Cs2SnSe3. The stronger bonding strength of Ge–Se is also evident from its larger force constant $\left|\Phi_{ij}\right|$ in comparison to that of Sn–Se (Figure S5). Hence, the shorter phonon lifetime of Cs2GeSe3 relative to Cs2SnSe3 may be attributed to their significant difference in Ge–Se and Sn–Se covalent bonding, which can be further quantified at the structural descriptor level through subsequent interpretable ML approaches.
@@ III Conclusion
In summary, we propose a novel HiBoFL framework via integrating unsupervised learning and supervised learning to efficiently predict complex properties and uncover structure-property relationships. As a compelling demonstration, this framework has been applied to efficiently identify semiconductors with ultralow $\kappa_{\mathrm{L}}$, which circumvents large-scale brute-force ab initio calculations without clear objectives. By employing unsupervised learning for the materials from the MP database, we successfully group them into seven clusters. A few hundred materials in clusters C1 and C2 with a high likelihood of possessing low $\kappa_{\mathrm{L}}$ is selected from a pool of hundreds of thousands based on similarity design rules. We further conduct low-cost HTC on materials belonging to the two clusters, establishing a local database for researchers to retrieve and providing a list of candidate materials with potential applications in the TE field. Additionally, statistical analysis of these candidates offers valuable domain knowledge to guide the design of materials with ultralow $\kappa_{\mathrm{L}}$. Cs2GeSe3 and Cs2SnSe3 with ultralow $\kappa_{\mathrm{L}}$ ($\sim$0.25 W/mK) are screened out, in which the anomalous ultralow $\kappa_{\mathrm{L}}$ of Cs2GeSe3 is attributed to the lower $\tau$ caused by the difference in covalent bonding. Based on the established local database, we train a robust ML classification model to refine unsupervised learning, achieving a ROC AUC of 0.94 on the test set, which can enable the efficient predictions of ultralow $\kappa_{\mathrm{L}}$ materials. The interpretable analysis of the classification model reveals the mechanisms of key descriptors on $\kappa_{\mathrm{L}}$ modulation, such as $W_{\mathrm{A}}^{\mathrm{min}}$, $\alpha$ and $L^{\mathrm{min}}$, etc. The special factor $L^{\mathrm{min}}$ is discovered to show a unique influence on structural anharmonicity, leading to the difference of phonon lifetime in Cs2GeSe3 and Cs2SnSe3, in agreement with the first-principles analysis. We believe that this work provides a novel feasible way for efficiently seeking promising TE materials, in which the proposed HiBoFL framework is also expected to be applied in other material fields.
@@ IV.2 Theoretical Framework
Where $\overline{\upsilon}_{\mathrm{l}}$, $\overline{\upsilon}_{\mathrm{t}}$, $\overline{\gamma}_{\mathrm{l}}$ and $\overline{\gamma}_{\mathrm{t}}$ are the longitudinal sound velocity, transverse sound velocity, longitudinal Grüneisen parameter, and transverse Grüneisen parameter, respectively. B and G are bulk moduli and shear moduli, respectively. In this manner, we can calculate the $\kappa_{\mathrm{PET}}$ for different materials derived from their elastic properties within the HTC framework.
To further validate $\kappa_{\mathrm{PET}}$ from the PET empirical equation and investigate the phonon thermal transport mechanisms, we calculated more accurate $\kappa_{\mathrm{L}}$ values for Cs2SnSe3 and Cs2SnSe3 by iteratively solving the phonon BTE:[88, 91]
$\kappa_{\mathrm{L}}^{\alpha\beta}=\frac{1}{k_{\mathrm{B}}T^{2}\Omega N}\sum_{\lambda}f_{0}\left(f_{0}+1\right)\left(\hbar\omega_{\lambda}\right)^{2}\upsilon_{\lambda}^{\alpha}F_{\lambda}^{\beta}$ (8)
Where $\alpha$ and $\beta$ are the Cartesian indexes. $k_{\mathrm{B}}$, $T$, $\Omega$, $N$ are the Boltzmann constant, temperature, volume of the unit cell, and regular grid of q points, respectively. $\lambda$ is a phonon mode including the branch index p and wave vector q, and $f_{0}$ is the phonon distribution function based on Bose-Einstein statistics. $\hbar$, $\omega_{\lambda}$, $\upsilon_{\lambda}^{\alpha}$ are the reduced Planck constant, phonon frequency, and phonon group velocity along the $\alpha$ direction, respectively. When only considering two- and three-phonon processes that contribute to scattering, the linearized BTE $F_{\lambda}$ takes the form:
$F_{\lambda}=\tau_{\lambda}^{0}(\upsilon_{\lambda}+\Delta_{\lambda})$ (9)
Where $\tau_{\lambda}^{0}$ and $\Delta_{\lambda}$ are the phonon lifetime of mode $\lambda$ and corrective term from iteration, respectively.
IV.3 Machine Learning
The chemical compositions and crystal structures of materials were featurized into different descriptors using the Matminer package.[92] All parts related to ML were carried out using the Scikit-learn package.[93] To achieve the automation and acceleration of hyperparameter optimization in supervised learning, we employed the powerful Optuna package[94] for efficiently finding the best hyperparameters. And to provide the interpretable analysis of the black-box ML model, a game-theoretic approach as implemented in the SHAP package[48, 49, 50] was employed.
For unsupervised learning, the elbow method and the silhouette coefficient method are utilized to determine the optimal k value in k-means clustering. In the elbow method, the sum of squared errors (SSE) is defined as:
$SSE=\sum_{i=1}^{k}\sum_{x\in C_{i}}\|x-\mu_{i}\|^{2}$ (10)
Where $C_{i}$ represents the set of data points in cluster i, $x$ is a data point in $C_{i}$, $\mu_{i}$ is the centroid of $C_{i}$. The silhouette coefficient is defined as:
$s(x)=\frac{b(x)-a(x)}{\max(a(x),b(x))}$ (11)
Where $b(x)$ is the average intra-cluster distance and measures how close the data point x is to other points in $C_{i}$, $a(x)$ is the average nearest-cluster distance and measures how far data point x is from the closest other cluster. After calculating the silhouette coefficient for all data points, the average value is then taken to obtain the mean silhouette coefficient for different values of k.
For supervised learning, accuracy, precision, recall, F-score, and ROC AUC are utilized to evaluate the performance metrics of classification models, which are defined as:
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (12)
$Precision=\frac{TP}{TP+FP}$ (13)
$Recall=\frac{TP}{TP+FN}$ (14)
$F\text{-}socre=\frac{2TP}{2TP+FP+FN}$ (15)
$ROC\ AUC=\int_{0}^{1}TPR\,d(FPR)$ (16)
Where TP is the number of true positives, FP is the number of false positives, TN is the number of true negatives, and FN is the number of false negatives. Moreover, TPR and FPR are the true positive rate and false positive rate, respectively, which are defined as:
$TPR=\frac{TP}{TP+FN}$ (17)
$FPR=\frac{FP}{TN+FP}$ (18)
For interpretable analysis, the SHAP value of a material descriptor is given by:
$\varphi_{i}(p)=\sum_{S\subseteq F\setminus\{i\}}\frac{|S|!(|F|-|S|-1)!}{|F|!}\left(p(S\cup\{i\})-p(S)\right)$ (19)
Where $p$ is the prediction model, $F$ represents the set of all material descriptors, $S$ represents the subset of all descriptors excluding the feature i, $p(S)$ is the model prediction when only descriptors in $S$ are considered, and $p(S\cup\{i\})$ is the model prediction when feature i is also included.